Find the value of : 3 tan2 450 - sin2
600 - cos2
300 - sec2
450

Answer by ikleyn(52855)

.
Find the value of : 3 tan2 450 - sin2
600 - cos2
300 - sec2
450
~~~~~~~~~~~~~~~~~`

tan(450) = tan(450-360) = tan(90) = INFINITY.


Computer says NAN in this case, which means "this number is not defined".